OpenStudy (anonymous):

solve the equation x=3b+2y A)y=2x+3b B)y=x-3b/2 C)y=x+3b/2 D)y=2(x+b)

OpenStudy (zepp):

Here, we are asking you to put y on one side.

OpenStudy (zepp):

Can you move the 2y from the right side to the left side?

OpenStudy (anonymous):

y= (x-3b/2)

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

subtract 3b frm both sides x-3b =2y then divide by 2

OpenStudy (zepp):

x=3b+2y You have this right? If you moved 3b to the left side you'll get (sorry I made a mistake above :)) x-3b = 2y Then divide by 2 to take off the 2 (x-3b)/2 = y

OpenStudy (anonymous):

i don't understand divide 2 to what?